OpenStudy (anonymous):

5. Solve this system x=2y-8 4x+y=13 A. (-2,-5) B. (2,-5) C. (2,5) D. (-2,5)

OpenStudy (texaschic101):

x = 2y - 8 now sub 2y - 8 in for x in the other equation 4x + y = 13 4(2y - 8) + y = 13 -- distribute through the parenthesis 8y - 32 + y = 13 8y + y = 13 + 32 9y = 45 y = 5 now sub 5 in for y in either equation to find x x = 2y - 8 x = 2(5) - 8 x = 10 - 8 x = 2 solution is : (2,5)

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

sure thing :)